# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v1

@@ C.2 Set-based MI Verification
For completeness, we evaluate how effective their approach is, if the users are able to mislabel their data for provenance purpose. We follow the targeted label flipping attack in Tramer et al., where for each target instance $(x,y)$, we inject multiple mislabeled data points $D_{\text{adv}}=\{(x,y^{\prime}),...,(x,y^{\prime})\}$ for some label $y^{\prime}\neq y$ [65]. We inject different number of poisoned samples and compare their approach with MembershipTracker.
@@
Table 10 shows the results, where the method by Tramer et al. indeed increases the exposure against the targeted samples, and the FPR@100% TPR is reduced from 49.33$\sim$56.32to 4.92$\sim$10.80%. Fig. 16 shows the ROC curve under the setting of 15 samples per user (similar trend on other settings). Meanwhile, the second row in Table 10 also shows that injecting more poisoned samples can contribute to higher MI success, which is also in line with [65]. Nevertheless, even in the case where each user contributes 25 samples, the method by Tramer et al. still incurs 4.92% FPR@100% TPR; while MembershipTracker has 0% FPR@100% TPR.
@@
Table 10: Comparing the technique by Tramer et al. and MembershipTracker in terms of their MI effectiveness for tracing data provenance.
@@
Figure 16: Comparing the performance between the approach by Tramer et al. and MembershipTracker (under 15 samples per user), and MembershipTracker achieves much higher MI success than their approach.
Appendix F Evaluating Huang et al. [31] under Limited Data Setting
Concurrent work by Huang et al. introduces a general framework for auditing unauthorized data use in training DL models [31]. Their main idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published Vs. the unpublished one.
@@
To understand the performance of their technique under such a limited-data setting, we use the original implementation from the authors [31] for an evaluation. Their approach can be applied to both image classifiers and foundation models, and we evaluate their approach under the same classifier setting as MembershipTracker (using CIFAR100). We first consider marking 1% of the dataset, and find that their technique can still maintain 100% detection success rate.
We then stress test their technique using 0.2% marking percentage, under which their auditing performance degrades considerably: the detection success rate is reduced to 65% when the output vector is available. Under the label-only setting, the success rate further drops to 25%. Therefore, our evaluation indicates that auditing data use under the limited-data setting still remains a major challenge, and MembershipTracker represents a desirable technique that can offer superior auditing performance in these settings.
@@
Appendix G Detailed Results for Output Perturbation Defense
Section 5.2.7 studies two defense strategies to perturb the model’s outputs: (1) injecting Gaussian noise to the prediction outputs; and (2) returning only the top-1 prediction label. We discuss the detailed results below.
Figure 17: Output noise injection increases the prediction loss of the member samples (from the leftmost to the middle figure), but their loss are still lower than many non-member samples’. Thus, MembershipTracker can leverage the average loss of the samples by each user to achieve a 0% FPR@ 100% TPR (the rightmost figure).

### v2

@@ C.2. Ablation of the Set-based MI Verification
For completeness, we evaluate how effective their approach is, if the users are able to mislabel their data for provenance purpose. We follow the targeted label flipping attack in Tramer et al., where for each target instance $(x,y)$, we inject multiple mislabeled data points $D_{\text{adv}}=\{(x,y^{\prime}),...,(x,y^{\prime})\}$ for some label $y^{\prime}\neq y$ (Tramèr et al., 2022). We inject different number of poisoned samples and compare their approach with MembershipTracker.
@@
Table 10 shows the results, where the method by Tramer et al. indeed increases the exposure against the targeted samples, and the FPR@100% TPR is reduced from 49.33$\sim$56.32to 4.92$\sim$10.80%. Fig. 17 shows the ROC curve under the setting of 15 samples per user (similar trend on other settings). Meanwhile, the second row in Table 10 also shows that injecting more poisoned samples can contribute to higher MI success, which is also in line with (Tramèr et al., 2022). Nevertheless, even in the case where each user contributes 25 samples, the method by Tramer et al. still incurs 4.92% FPR@100% TPR; while MembershipTracker has 0% FPR@100% TPR.
@@
Table 10. Comparing the technique by Tramer et al. and MembershipTracker in terms of their MI effectiveness for tracing data provenance.
@@
Figure 17. Comparing the performance between the approach by Tramer et al. and MembershipTracker (under 15 samples per user), and MembershipTracker achieves much higher MI success than their approach.
Appendix F Evaluating Huang et al. (Huang et al., 2024) under Limited Data Setting
Recent work by Huang et al. introduces a general framework for auditing unauthorized data use in training DL models (Huang et al., 2024). Their main idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published Vs. the unpublished one.
@@
To understand the performance of their technique under such a limited-data setting, we use the original implementation from the authors (Huang et al., 2024) for an evaluation. Their approach can be applied to both image classifiers and foundation models, and we evaluate their approach under the same classifier setting as MembershipTracker (using CIFAR100). We first consider marking 1% of the dataset, and find that their technique can still maintain 100% detection success rate.
We then stress test their technique using 0.1% marking percentage, under which their auditing performance degrades considerably: the detection success rate is reduced to 20% when the output vector is available. Under the label-only setting, the success rate further drops to 5%. Therefore, our evaluation indicates that auditing data use under the limited-data setting still remains a major challenge, and MembershipTracker represents a desirable technique that can offer superior auditing performance in these settings.